1. Possible Lewis structures for 6 molecules are shown below. Which figures show correct structures? (Use the drop down menusPossible Lewis structures for 6 molecules are shown below. Which figures show correct structures? (Use the drop down menus to choose "Correct" or "Incorrect" for each structure.
Answer #1

The Lewis structures show all the valance electrons of the bonded atoms. Now the shared pairs of the valance electrons make a bond (represented by a line) and the unshared pairs remain as lone pair electrons, (represented by a pair of dots) to complete their octet.

The octet means total 8 electrons around an atom

Now total electron = bond pair electrons + lone pair electrons

= (total bond of an atom in the structure * 2) + (total lone pair of that atom is structure * 2)

Thus each single line or bond will be calculated as 2 electrons.

Now in the given structures-

1-H- 6 =.5-H e of e=4 Lue know valance H=1 s-6 How in the given stonetune. ( The band pair of I will be calculated HRB sooth fototal 9=2 tot 4: total 2:8 ( (4: total - x tatal suo troetel eslo More than 8 e re Thus, song sto] 6:00 ⓇHODOD-22 Since all ttotal g@ Octet Rule sure Herce all atoms correct salisily spel - (sh) (H 29 . (6+2=80) Sire all the aterm follow Oct et ruler